# You May Use the Same Channel Knowledge Map for Environment-Aware NLoS Sensing and Communication

### 17 Mar 2026 v3

@@ Abstract
As one of the key usage scenarios for the sixth generation (6G) wireless networks, integrated sensing and communication (ISAC) provides an efficient framework to achieve simultaneous wireless sensing and communication. However, traditional wireless sensing techniques mainly rely on the line-of-sight (LoS) assumptions, i.e., the sensing targets are directly visible to both the sensing transmitter and receiver. This hinders ISAC systems to be applied in complex environments such as the urban low-altitude airspace, which usually suffers from signal blockage and non-line-of-sight (NLoS) multi-path propagation. To address this challenge, in this paper, we propose a novel approach to enable environment-aware NLoS ISAC by leveraging the new technique called channel knowledge map (CKM), which was originally proposed for environment-aware wireless communications. One major novelty of our proposed method is that the same CKM built for wireless communication can be directly used to enable NLoS wireless sensing, thus enjoying the benefits of “killing two birds with one stone”. To this end, the sensing targets are treated as virtual user equipment (UE), and the wireless communication channel priors are transformed into the sensing channel priors, allowing one single CKM to serve dual purposes. We illustrate our proposed framework by a specific CKM called channel angle-delay map (CADM). Specifically, the proposed framework utilizes CADM to derive angle-delay priors of the sensing channel by exploiting the relationship between communication and sensing angle-delay distributions, enabling sensing target localization in the challenging NLoS environment. Extensive simulation results demonstrate significant performance improvements over classic geometry-based sensing methods, which is further validated by Cramér-Rao Lower Bound (CRLB) analysis.
@@ I Introduction
Significant progress has been made for ISAC systems, particularly in line-of-sight (LoS) scenarios, where the sensing targets are directly visible to both the transmitter and receiver. Various ISAC techniques have been developed, like joint waveform design, beamforming optimization, and interference management to balance radar-like sensing with data transmission, achieving high localization accuracy and spectral efficiency in LoS conditions [37, 29, 15, 46, 49]. For instance, methods such as time-of-arrival (ToA) estimation and angle-based localization leverage direct LoS paths to estimate target locations [36, 25]. However, in complex environments such as urban low-altitude airspace, LoS paths are frequently blocked. As a result, sensing in such environments can only rely on pure non-line-of-sight (NLoS) signal paths, posing a much greater challenge to map the estimated sensing parameters to the location of the sensing targets as more unknowns are involved [21, 26]. Some techniques have been proposed to achieve NLoS localization, including fingerprinting [31], single-base-station (BS) localization exploiting near-field scatterers [48], and reconfigurable intelligent surface (RIS)-aided methods [16]. However, they cannot be applied for sensing since unlike localization tasks, the sensing targets are typically non-cooperative in the sense that they only passively reflect the signals rather than transmitting or receiving them. Specifically, fingerprinting based localization works by building a database that maps user equipment (UE) location to its wireless communication channel signature such as received signal strength (RSS). However, this approach is unsuitable for sensing. The fundamental reason is that a fingerprint database characterizes the direct channel between UE and the anchor nodes, whereas sensing relies on the echoes from the passive sensing target. Therefore, the fingerprint database built for localization cannot be used for wireless sensing. Instead, sensing would require its own “sensing database” that fingerprints target-specific echo characteristics. Creating such databases is quite challenging, if not impossible, as different types of targets (e.g., vehicles, pedestrians) would need a distinct and comprehensive map of its potential echo signatures. On the other hand, single-BS near-field localization studied in [48] first senses near-field scatterers to acquire environment priors, and then leverages the location of near-field scatterers to localize UEs without relying on LoS link. However, it relies on near-field assumptions, limiting applicability in far-field ISAC. In addition, RIS-aided methods create virtual LoS relationships for localization by deploying reflective surfaces, they require additional costly and extensive hardware deployments. In fact, rigorous theoretical studies have proven that, without environment or NLoS prior knowledge, NLoS paths cannot provide beneficial information for localization [24]. This demonstrates the imperative need to develop an innovative framework that can leverage environment prior knowledge to tackle the challenging problem of NLoS sensing for ISAC systems.
@@
Channel knowledge map (CKM) is a promising technique to achieve the above goal [43, 42]. CKM is a site-specific database capturing location-specific channel characteristics—such as angles of arrival (AoAs), angles of departure (AoDs), delays, and path gains—learned from historical UE-BS interactions. This definition distinguishes CKM from the conventional radio map. While radio maps are invaluable for many applications, such as fingerprinting-based localization or interference management [10], they usually store signal quality metrics (such as RSS) that are dependent on transmitter and receiver setup and activity. In contrast, CKM aims to directly reflect the intrinsic wireless channel properties of the local environment, independent of transmitter/receiver activity. As “knowledge” can take various forms, the term CKM aims to signify the much richer, multi-dimensional characteristics of the channel’s underlying physical parameters (e.g., AoAs, AoDs, and delays), which are essential for advanced applications like environment-aware communication and sensing. Recently, extensive research efforts have been devoted to CKM construction and CKM-based environment-aware communication and sensing. For example, to enable AI-based intelligent CKM construction, datasets like CKMImageNet have been developed [35]. Besides, theoretical analyses have quantified the data requirements for CKM construction [38], and various methods have been proposed for CKM construction, leveraging techniques such as generative AI models for sparsely observed data, path matching, and the fusion of environment information like point clouds [12, 8, 14, 32, 17]. In terms of CKM utilization for wireless communications, various problems have been investigated, including optimizing resource allocation [41, 18], improving interference management [11], enabling efficient beam alignment in mmWave systems [34, 4], facilitating RIS-aided communications [27], and enhancing channel prediction [30]. Besides, prototyping systems have been developed that validated its ability to achieve environment-aware mmWave beam alignment, improving connectivity in complex indoor settings [4]. Recent efforts have also explored CKM for sensing applications, such as environment-aware channel estimation with dynamic sensing data [33], or to enhance CKM construction and applications by sensing the target locations [44]. Besides, CKM is utilized to enhance sensing by suppressing environmental clutter, improving target parameter estimation [39]. These advancements have established a strong foundation for CKM-based sensing. However, they have primarily focused on leveraging CKM for sensing-assisted communication or sensing scenarios by assuming the existence of LoS links, leaving the full potential of exploiting NLoS paths in complex environments underexplored. This gap inspires the development of a novel ISAC framework that directly utilizes NLoS paths to extract target information, leveraging the environmental priors embedded in CKM to address NLoS challenges effectively.
@@
Therefore, in this paper, we try to answer the following question: Can we use the same CKM to achieve both environment-aware NLoS sensing and communication? We show that the answer is affirmative, by proposing a novel CKM-based environment-aware ISAC framework. Different from the conventional environment-unaware ISAC approaches, which rely on the existence of LoS path, the proposed method can exploit NLoS paths for sensing even when LoS path is absent. The main contributions of this work are summarized as follows:
@@
  • NLoS ISAC model: We propose a novel system model for ISAC in NLoS scenarios without relying on those strong assumptions, such as the presence of LoS path and reciprocal signal propagation where AoA equals AoD. Specifically, unlike existing models that usually assume signals return along the same path, failing to capture NLoS environments where signals may reflect off different scatterers, our model supports both reciprocal and non-reciprocal paths, allowing distinct AoA and AoD. This enables comprehensive path modeling, capturing all possible NLoS path combinations, enhancing applicability to diverse propagation scenarios and establishing a foundation for environment-aware NLoS ISAC.
@@
  • Fundamental principle of applying CKM to sensing: This work elucidates the core principle of utilizing CKM, originally developed for communication [43, 42], to enable environment-aware NLoS sensing, demonstrating that the same CKM can be used for both environment-aware NLoS sensing and communication. To this end, by treating sensing targets as virtual UE, CKM leverages location-specific channel knowledge (such as angles, delays) to infer sensing parameters. This is a non-trivial task, since a CKM built for communication learns one-way location-specific channel statistical priors, which cannot be directly applied to the two-way cascaded sensing channel. Our primary contribution is the novel mathematical transformation framework that bridges this gap. This approach utilizes environment priors embedded in CKM to exploit NLoS paths effectively, overcoming the limitations of environment-unaware methods that fail to map those estimated sensing parameters (such as multi-path angles and delays) to target locations in NLoS scenarios.
@@
  • Sensing based on channel angle-delay map (CADM): We propose a particular type of CKM, named CADM, which tries to learn location-specific angle and delay distributions of channel multipath. For each path, CADM uses certain distributions like Gaussian distribution to characterize the angle and delay. The parameters of this distribution—mean and variance—are dependent on the specific location. As this dependency is usually highly complex and non-linear, making analytical modeling difficult, we employ a fully connected neural network (FCNN) to learn the mapping from location to these statistical parameters. By treating sensing target as virtual UE, the communication channel knowledge distributions in the CADM, originally tied to UE location, are transformed into a sensing likelihood function corresponding to the sensing target’s location. Then, for target sensing, we perform maximum likelihood (ML) estimation. This process utilizes both the measured angle-delay data and the sensing likelihood function derived from CADM, and is optimized via gradient descent. The proposed CADM-based method achieves environment-aware sensing by utilizing NLoS path information, where traditional LoS-based methods fail in pure NLoS environments.
@@
  • Extensive evaluation and superior performance: Extensive simulations in the challenging NLoS environments validate the proposed method, demonstrating significant improvements in sensing accuracy compared to environment-unaware approaches, such as geometric triangulation. A rigorous performance analysis using the Cramér-Rao lower bound (CRLB) verified the benefits of multipath exploitation in environment-aware sensing enabled by CKM.
@@
The remainder of this paper is organized as follows. Section II presents the system model for the ISAC framework. Section III discusses the motivation for CKM-enabled sensing, while Section IV details the mechanisms for the proposed CADM-enabled NLoS sensing, respectively. Section V provides simulation results, and Section VI concludes the paper.
@@ II System Model
Consider an ISAC system, as shown in Fig. 1, where a BS located at $\mathbf{b}=(x_{\mathbf{b}},y_{\mathbf{b}})$ tries to communicate with a single-antenna UE and simultaneously provide sensing services. The BS is equipped with $N$ antennas for signal transmission and $N_{\text{Rx}}$ antennas for sensing signal reception. The UE is located at $\mathbf{q}=(x_{\mathbf{q}},y_{\mathbf{q}})$. While communicating with the UE, ISAC system tries to estimate the location $\mathbf{x}=(x_{\mathbf{x}},y_{\mathbf{x}})$ of the sensing object (e.g., vehicles or pedestrians) in the environment. We consider the mono-static sensing mode where the BS operates in full-duplex, utilizing the $N$ antennas for transmitting signals and the $N_{\text{Rx}}$ receive antennas to process the received echoes. Different from the conventional ISAC systems [3], we focus on the challenging NLoS scenario where the LoS path between the BS and the target is blocked, relying solely on NLoS paths for target sensing, as illustrated in Fig. 1.
@@
where $\mathbf{s}(t)\in\mathbb{C}^{N\times 1}$ is the transmitted signal, $\alpha$ is the complex gain of the path, $\mathbf{a}_{r}(\theta)\in\mathbb{C}^{N_{\mathrm{Rx}}\times 1}$ and $\mathbf{a}_{t}(\theta)\in\mathbb{C}^{N\times 1}$ are the steering vectors corresponding to the AoA $\theta$ for the $N_{\text{Rx}}$ receive antennas and the AoD $\theta$ for the $N$ transmit antennas, respectively. Besides, $\tau$ is the round-trip delay and $v$ is the Doppler shift, $\mathbf{n}\sim\mathcal{CN}(\mathbf{0},\sigma^{2}\mathbf{I}_{N_{\text{Rx}}})$ is the zero-mean circularly symmetric complex Gaussian (CSCG) noise vector with variance $\sigma^{2}$. Note that to gain the essential insights, we only focus on the echo signal from the sensing target by assuming that clutter signal has been properly removed [21]. In such a LoS scenario, it is further assumed for geometric consistency that the BS is equipped with a uniform linear array (ULA) oriented along the y-axis as shown in Fig. 2. Therefore, the AoA/AoD $\theta$ and delay $\tau$ are directly related to the locations of the BS $\mathbf{b}$ and the sensing target $\mathbf{x}$, expressed as
@@
However, the above result is only applicable when the LoS link between the BS and sensing targets exists. In the challenging NLoS scenario where the LoS link is blocked, as illustrated in Fig. 2(b), the signal transmitted by the BS arrives the sensing target via forward paths with reflections, diffractions, or scattering, echoed by the sensing target and then returns to the BS via reverse paths. In this case, there exist multiple paths between the BS and the target, each characterized by distinct AoA, AoD, delays, and Doppler shifts. The received echo signal at the BS, denoted as $\mathbf{y}(t)$, is thus expressed as
@@
where $\alpha_{l_{T},l_{R}},\tau_{l_{T},l_{R}}$ and $\nu_{l_{T},l_{R}}$ respectively denote the complex gain, delay, and Doppler shift for the path combination $(l_{T},l_{R})$. This formula includes all the $L=L^{\prime 2}$ combinations of forward and reverse paths.
@@
To ultimately achieve the sensing goal, these parameters in $\mathbf{z}$ must be mapped to the location $\mathbf{x}$ of the sensing target. In LoS scenarios, this mapping can be directly achieved with the closed-form expression (3). However, such a mapping is no longer valid for the challenging scenario where LoS is absent, since NLoS paths dependent on the environment, denoted as $E$, as shown in Fig. 1. Therefore, when only NLoS paths are available, without any prior knowledge about the environment, determining the sensing target location $\mathbf{x}$ based on the sensed channel parameters $\mathbf{z}$ is an underdetermined problem. In fact, even in mixed scenarios with both LoS and NLoS paths, conventional methods typically employ techniques such as NLoS mitigation or path selection to isolate the LoS path, discarding NLoS paths that contain valuable target-related information [20]. This results in information loss, limiting the sensing accuracy. To utilize the target-related information in NLoS paths, which depends on $E$, an environment-aware sensing model that incorporates $E$ is necessary to leverage both LoS (if it exists) and NLoS paths for sensing target localization.
@@
where $f(\cdot)$ is an abstract representation of the mapping from the sensing target location $\mathbf{x}$ and environment $E$ to the observed channel knowledge, denoted as $\hat{\mathbf{z}}$. In the LoS case, where $E=\emptyset$, the mapping $f(\cdot)$ degenerates to the closed-form expression in (2). The vector $\mathbf{e}\in\mathbb{R}^{3L\times 1}$ models the errors associated with the angle and delay estimations, assumed to follow independent multivariate Gaussian distribution $\mathbf{e}\sim\mathcal{N}(\mathbf{0},\mathbf{\Sigma}_{e})$, with
@@
where $P(\hat{\mathbf{z}}|\mathbf{x})$ represents the likelihood function of the observed channel knowledge $\hat{\mathbf{z}}$ shown at the top of the next page. With Gaussian error $\mathbf{e}$, the log-likelihood function can be expressed as
$\ln P(\hat{\mathbf{z}}|\mathbf{x})=\mathcal{N}(\hat{\mathbf{z}};f(\mathbf{x},E),\mathbf{\Sigma}_{e}).$ (11)
The problem (10) can be solved by classic optimization methods such as gradient descent method. The update rule for the target location at the $(k+1)$th iteration is given by
@@
$\nabla_{\mathbf{x}}\left(-\ln P(\hat{\mathbf{z}}|\mathbf{x})\right)=\mathbf{J}^{T}\mathbf{\Sigma}_{e}^{-1}\left(\hat{\mathbf{z}}-f(\mathbf{x},E)\right),$ (13)
@@
However, the practical implementation of environment-aware sensing based on (12) is highly non-trivial. The first major difficulty lies in acquiring an accurate representation of the environment $E$ and formulating its mathematical model. The second challenge arises that even if $E$ is known, the mapping function $f(\cdot)$ remains difficult to determine precisely. Potential approaches, such as solving Maxwell’s equations or employing ray tracing techniques, suffer from prohibitively high computational complexity, rendering them unsuitable for real-time sensing applications. Furthermore, these approaches cannot guarantee differentiability, which is essential for gradient-based optimization methods in ML estimation as (12).

### 09 Apr 2026 v4

@@ Abstract
As one of the key usage scenarios for the sixth generation (6G) wireless networks, integrated sensing and communication (ISAC) provides an efficient framework to achieve simultaneous wireless sensing and communication. However, traditional wireless sensing techniques mainly rely on line-of-sight (LoS) assumptions, i.e., the sensing targets are directly visible to both the sensing transmitter and receiver. This prevents ISAC systems from being applied in complex environments such as the urban low-altitude airspace, which usually suffers from signal blockage and non-line-of-sight (NLoS) multipath propagation. To address this challenge, in this paper, we propose a novel approach to enable environment-aware NLoS ISAC by leveraging the new technique called channel knowledge map (CKM), which was originally proposed for environment-aware wireless communications. One major novelty of our proposed method is that the same CKM built for wireless communication can be directly used to enable NLoS wireless sensing, thus enjoying the benefits of “killing two birds with one stone”. To this end, the sensing targets are treated as virtual user equipment (UE), and the wireless communication channel priors are transformed into the sensing channel priors, allowing one single CKM to serve dual purposes. We illustrate our proposed framework using a specific CKM called channel angle-delay map (CADM). Specifically, the proposed framework utilizes CADM to derive angle-delay priors of the sensing channel by exploiting the relationship between communication and sensing angle-delay distributions, enabling sensing target localization in the challenging NLoS environment. Extensive simulation results demonstrate significant performance improvements over classic geometry-based sensing methods, which are further validated by Cramér-Rao Lower Bound (CRLB) analysis.
@@ I Introduction
Significant progress has been made in ISAC systems, particularly in line-of-sight (LoS) scenarios, where the sensing targets are directly visible to both the transmitter and receiver. Various ISAC techniques have been developed, like joint waveform design, beamforming optimization, and interference management to balance radar-like sensing with data transmission, achieving high localization accuracy and spectral efficiency in LoS conditions [37, 29, 15, 46, 49]. For instance, methods such as time-of-arrival (ToA) estimation and angle-based localization leverage direct LoS paths to estimate target locations [36, 25]. However, in complex environments such as urban low-altitude airspace, LoS paths are frequently blocked. As a result, sensing in such environments can only rely on pure non-line-of-sight (NLoS) signal paths, posing a much greater challenge to map the estimated sensing parameters to the location of the sensing targets as more unknowns are involved [21, 26]. Some techniques have been proposed to achieve NLoS localization, including fingerprinting [31], single-base-station (BS) localization exploiting near-field scatterers [48], and reconfigurable intelligent surface (RIS)-aided methods [16]. However, they cannot be applied to sensing, because unlike localization tasks, the sensing targets are typically non-cooperative in the sense that they only passively reflect the signals rather than transmitting or receiving them. Specifically, fingerprinting-based localization works by building a database that maps user equipment (UE) location to its wireless communication channel signature such as received signal strength (RSS). However, this approach is unsuitable for sensing. The fundamental reason is that a fingerprint database characterizes the direct channel between the UE and the anchor nodes, whereas sensing relies on the echoes from the passive sensing target. Therefore, the fingerprint database built for localization cannot be used for wireless sensing. Instead, sensing would require its own “sensing database” that fingerprints target-specific echo characteristics. Creating such databases is quite challenging, if not impossible, as different types of targets (e.g., vehicles, pedestrians) would need a distinct and comprehensive map of their potential echo signatures. On the other hand, single-BS near-field localization studied in [48] first senses near-field scatterers to acquire environment priors, and then leverages the location of near-field scatterers to localize UEs without relying on a LoS link. However, it relies on near-field assumptions, limiting applicability in far-field ISAC. In addition, while RIS-aided methods create virtual LoS relationships for localization by deploying reflective surfaces, they require additional costly and extensive hardware deployments. In fact, rigorous theoretical studies have proven that, without environment or NLoS prior knowledge, NLoS paths cannot provide beneficial information for localization [24]. This demonstrates the imperative need to develop an innovative framework that can leverage environment prior knowledge to tackle the challenging problem of NLoS sensing for ISAC systems.
@@
Channel knowledge map (CKM) is a promising technique to achieve the above goal [43, 42]. CKM is a site-specific database capturing location-specific channel characteristics—such as angles of arrival (AoAs), angles of departure (AoDs), delays, and path gains—learned from historical UE-BS interactions. This definition distinguishes CKM from the conventional radio map. While radio maps are invaluable for many applications, such as fingerprinting-based localization or interference management [10], they usually store signal quality metrics (such as RSS) that are dependent on transmitter and receiver setup and activity. In contrast, CKM aims to directly reflect the intrinsic wireless channel properties of the local environment, independent of transmitter/receiver activity. As “knowledge” can take various forms, the term CKM aims to signify the much richer, multi-dimensional characteristics of the channel’s underlying physical parameters (e.g., AoAs, AoDs, and delays), which are essential for advanced applications like environment-aware communication and sensing. Recently, extensive research efforts have been devoted to CKM construction and CKM-based environment-aware communication and sensing. For example, to enable AI-based intelligent CKM construction, datasets like CKMImageNet have been developed [35]. Besides, theoretical analyses have quantified the data requirements for CKM construction [38], and various methods have been proposed for CKM construction, leveraging techniques such as generative AI models for sparsely observed data, path matching, and the fusion of environment information like point clouds [12, 8, 14, 32, 17]. In terms of CKM utilization for wireless communications, various problems have been investigated, including optimizing resource allocation [41, 18], improving interference management [11], enabling efficient beam alignment in mmWave systems [34, 4], facilitating RIS-aided communications [27], and enhancing channel prediction [30]. Moreover, prototyping systems have been developed that have validated its ability to achieve environment-aware mmWave beam alignment, improving connectivity in complex indoor settings [4]. Recent efforts have also explored CKM for sensing applications, such as environment-aware channel estimation with dynamic sensing data [33], or to enhance CKM construction and applications by sensing the target locations [44]. Furthermore, CKM is utilized to enhance sensing by suppressing environmental clutter, improving target parameter estimation [39]. These advancements have established a strong foundation for CKM-based sensing. However, they have primarily focused on leveraging CKM for sensing-assisted communication or sensing scenarios by assuming the existence of LoS links, leaving the full potential of exploiting NLoS paths in complex environments underexplored. This gap inspires the development of a novel ISAC framework that directly utilizes NLoS paths to extract target information, leveraging the environmental priors embedded in CKM to address NLoS challenges effectively.
@@
Therefore, in this paper, we try to answer the following question: Can we use the same CKM to achieve both environment-aware NLoS sensing and communication? We show that the answer is affirmative, by proposing a novel CKM-based environment-aware ISAC framework. Different from the conventional environment-unaware ISAC approaches, which rely on the existence of a LoS path, the proposed method can exploit NLoS paths for sensing even when the LoS path is absent. The main contributions of this work are summarized as follows:
@@
  • NLoS ISAC model: We propose a novel system model for ISAC in NLoS scenarios without relying on those strong assumptions, such as the presence of a LoS path and reciprocal signal propagation where AoA equals AoD. Specifically, unlike existing models that usually assume signals return along the same path, failing to capture NLoS environments where signals may reflect off different scatterers, our model supports both reciprocal and non-reciprocal paths, allowing distinct AoA and AoD. This enables comprehensive path modeling, capturing all possible NLoS path combinations, enhancing applicability to diverse propagation scenarios and establishing a foundation for environment-aware NLoS ISAC.
@@
  • Fundamental principle of applying CKM to sensing: This work elucidates the core principle of utilizing CKM, originally developed for communication [43, 42], to enable environment-aware NLoS sensing, demonstrating that the same CKM can be used for both environment-aware NLoS sensing and communication. To this end, by treating sensing targets as virtual UE, CKM leverages location-specific channel knowledge (such as angles, delays) to infer sensing parameters. This is a non-trivial task, since a CKM built for communication learns one-way location-specific channel statistical priors, which cannot be directly applied to the two-way cascaded sensing channel. Our primary contribution is the novel mathematical transformation framework that bridges this gap. This approach utilizes environment priors embedded in CKM to exploit NLoS paths effectively, overcoming the limitations of environment-unaware methods that fail to map those estimated sensing parameters (such as multipath angles and delays) to target locations in NLoS scenarios.
@@
  • Sensing based on channel angle-delay map (CADM): We propose a particular type of CKM, named CADM, which tries to learn location-specific angle and delay distributions of channel multipath. For each path, CADM uses certain distributions like Gaussian distribution to characterize the angle and delay. The parameters of this distribution—mean and variance—are dependent on the specific location. As this dependency is usually highly complex and non-linear, making analytical modeling difficult, we employ a fully connected neural network (FCNN) to learn the mapping from location to these statistical parameters. By treating the sensing target as a virtual UE, the communication channel knowledge distributions in the CADM, originally tied to UE location, are transformed into a sensing likelihood function corresponding to the sensing target’s location. Then, for target sensing, we perform maximum likelihood (ML) estimation. This process utilizes both the measured angle-delay data and the sensing likelihood function derived from CADM, and is optimized via gradient descent. The proposed CADM-based method achieves environment-aware sensing by utilizing NLoS path information, where traditional LoS-based methods fail in pure NLoS environments.
@@
  • Extensive evaluation and superior performance: Extensive simulations in the challenging NLoS environments validate the proposed method, demonstrating significant improvements in sensing accuracy compared to environment-unaware approaches, such as geometric triangulation. A rigorous performance analysis using the Cramér-Rao lower bound (CRLB) verifies the benefits of multipath exploitation in environment-aware sensing enabled by CKM.
@@
The remainder of this paper is organized as follows. Section II presents the system model for the ISAC framework. Section III discusses the motivation for CKM-enabled sensing, while Section IV details the mechanisms for the proposed CADM-enabled NLoS sensing. Section V provides simulation results, and Section VI concludes the paper.
@@ II System Model
Consider an ISAC system, as shown in Fig. 1, where a BS located at $\mathbf{b}=(x_{\mathbf{b}},y_{\mathbf{b}})$ tries to communicate with a single-antenna UE and simultaneously provide sensing services. The BS is equipped with NN antennas for signal transmission and $N_{\text{Rx}}$ antennas for sensing signal reception. The UE is located at $\mathbf{q}=(x_{\mathbf{q}},y_{\mathbf{q}})$. While communicating with the UE, the ISAC system tries to estimate the location $\mathbf{x}=(x_{\mathbf{x}},y_{\mathbf{x}})$ of the sensing object (e.g., vehicles or pedestrians) in the environment. We consider the mono-static sensing mode where the BS operates in full-duplex, utilizing the NN antennas for transmitting signals and the $N_{\text{Rx}}$ receive antennas to process the received echoes. Different from the conventional ISAC systems [3], we focus on the challenging NLoS scenario where the LoS path between the BS and the target is blocked, relying solely on NLoS paths for target sensing, as illustrated in Fig. 1.
@@
where $\mathbf{s}(t)\in\mathbb{C}^{N\times 1}$ is the transmitted signal, $\alpha$ is the complex gain of the path, $\mathbf{a}_{r}(\theta)\in\mathbb{C}^{N_{\mathrm{Rx}}\times 1}$ and $\mathbf{a}_{t}(\theta)\in\mathbb{C}^{N\times 1}$ are the steering vectors corresponding to the AoA $\theta$ for the $N_{\text{Rx}}$ receive antennas and the AoD $\theta$ for the NN transmit antennas, respectively. Besides, $\tau$ is the round-trip delay and $v$ is the Doppler shift, $\mathbf{n}\sim\mathcal{CN}(\mathbf{0},\sigma^{2}\mathbf{I}_{N_{\text{Rx}}})$ is the zero-mean circularly symmetric complex Gaussian (CSCG) noise vector with variance $\sigma^{2}$. Note that to gain the essential insights, we only focus on the echo signal from the sensing target by assuming that the clutter signal has been properly removed [21]. In such a LoS scenario, it is further assumed for geometric consistency that the BS is equipped with a uniform linear array (ULA) oriented along the y-axis as shown in Fig. 2. Therefore, the AoA/AoD $\theta$ and delay $\tau$ are directly related to the locations of the BS $\mathbf{b}$ and the sensing target $\mathbf{x}$, expressed as
@@
However, the above result is only applicable when the LoS link between the BS and sensing targets exists. In the challenging NLoS scenario where the LoS link is blocked, as illustrated in Fig. 2(b), the signal transmitted by the BS arrives at the sensing target via forward paths with reflections, diffractions, or scattering, is reflected by the sensing target, and then returns to the BS via reverse paths. In this case, there exist multiple paths between the BS and the target, each characterized by distinct AoA, AoD, delays, and Doppler shifts. The received echo signal at the BS, denoted as $\mathbf{y}(t)$, is thus expressed as
@@
where $\alpha_{l_{T},l_{R}},\tau_{l_{T},l_{R}}$ and $v_{l_{T},l_{R}}$ respectively denote the complex gain, delay, and Doppler shift for the path combination $(l_{T},l_{R})$. This formula includes all the $L=L^{\prime 2}$ combinations of forward and reverse paths.
@@
To ultimately achieve the sensing goal, these parameters in $\mathbf{z}$ must be mapped to the location $\mathbf{x}$ of the sensing target. In LoS scenarios, this mapping can be directly achieved with the closed-form expression (3). However, such a mapping is no longer valid for the challenging scenario where LoS is absent, since NLoS paths depend on the environment, denoted as $E$, as shown in Fig. 1. Therefore, when only NLoS paths are available, without any prior knowledge about the environment, determining the sensing target location $\mathbf{x}$ based on the sensed channel parameters $\mathbf{z}$ is an underdetermined problem. In fact, even in mixed scenarios with both LoS and NLoS paths, conventional methods typically employ techniques such as NLoS mitigation or path selection to isolate the LoS path, discarding NLoS paths that contain valuable target-related information [20]. This results in information loss, limiting the sensing accuracy. To utilize the target-related information in NLoS paths, which depends on $E$, an environment-aware sensing model that incorporates $E$ is necessary to leverage both LoS (if it exists) and NLoS paths for sensing target localization.
@@
where $f(\cdot)$ is an abstract representation of the mapping from the sensing target location $\mathbf{x}$ and environment $E$ to the observed channel knowledge, denoted as $\hat{\mathbf{z}}$. In the LoS case, where $E=\emptyset$, the mapping $f(\cdot)$ degenerates to the closed-form expression in (2). The vector $\mathbf{e}\in\mathbb{R}^{3L\times 1}$ models the errors associated with the angle and delay estimations, assumed to follow an independent multivariate Gaussian distribution $\mathbf{e}\sim\mathcal{N}(\mathbf{0},\mathbf{\Sigma}_{e})$, with
@@
where $P(\hat{\mathbf{z}}|\mathbf{x})$ represents the likelihood function of the observed channel knowledge $\hat{\mathbf{z}}$. With Gaussian error $\mathbf{e}$, the log-likelihood function can be expressed as
$\ln P(\hat{\mathbf{z}}|\mathbf{x})=\ln\mathcal{N}(\hat{\mathbf{z}};f(\mathbf{x},E),\mathbf{\Sigma}_{e}).$ (11)
The problem (10) can be solved by classic optimization methods such as the gradient descent method. The update rule for the target location at the $(k+1)$th iteration is given by
@@
$\nabla_{\mathbf{x}}\left(-\ln P(\hat{\mathbf{z}}|\mathbf{x})\right)=-\mathbf{J}^{T}\mathbf{\Sigma}_{e}^{-1}\left(\hat{\mathbf{z}}-f(\mathbf{x},E)\right),$ (13)
@@
However, the practical implementation of environment-aware sensing based on (12) is highly non-trivial. The first major difficulty lies in acquiring an accurate representation of the environment $E$ and formulating its mathematical model. The second challenge is that even if $E$ is known, the mapping function $f(\cdot)$ remains difficult to determine precisely. Potential approaches, such as solving Maxwell’s equations or employing ray tracing techniques, suffer from prohibitively high computational complexity, rendering them unsuitable for real-time sensing applications. Furthermore, these approaches cannot guarantee differentiability, which is essential for gradient-based optimization methods in ML estimation as in (12).